OpenStudy (juan1857):

Write the equation in polar form. 3x-4y-10=0

OpenStudy (bahrom7893):

Do you know what the polar form of an equation is?

OpenStudy (juan1857):

yes I do

OpenStudy (bahrom7893):

Post it, what have you tried so far?

OpenStudy (juan1857):

when I found out that p=2, and currently struggling to get \[\phi\]

OpenStudy (juan1857):

as of right now I got Arctan 3/-4

OpenStudy (hyperpiper):

Oh My I cant do this sorry XDD

OpenStudy (juan1857):

oh, Its ok

OpenStudy (bahrom7893):

Okay, so: \[3x-4y-10=0\]\[3x-4y=10\]\[3r Cos\theta -4r Sin\theta = 10\] Now just solve for r and you're done.

OpenStudy (juan1857):

oh wow, I was taught a different way using arctan

OpenStudy (bahrom7893):

\[r(3Cos\theta - 4Sin\theta)=10\]\[r=\frac{10}{3Cos\theta-4Sin\theta}\]

OpenStudy (bahrom7893):

Hmm I have been out of college for a while, so I don't remember what I was taught - I was following this link to solve this though: http://www.ck12.org/book/CK-12-Trigonometry-Concepts/section/6.6/ (look at example 3)

OpenStudy (juan1857):

ok, thanks anyway

OpenStudy (kittiwitti1):

BACON CONFETTI!

OpenStudy (juan1857):

lol

OpenStudy (kittiwitti1):

remember what x and y in polar form are? \(x=r\cos{\theta}\text{ and }y=r\sin{\theta}\)

OpenStudy (juan1857):

yep

OpenStudy (kittiwitti1):

Okay so you have the equation \(3x-4y-10=0\) which becomes \(3r\cos{\theta}-4r\sin{\theta}-10=0\)

OpenStudy (kittiwitti1):

... I don't know what I'm doing ; - ;

OpenStudy (kittiwitti1):

I'm studying for a test that involves these things too and my brain's fried x_x

OpenStudy (juan1857):

just outta curiosity, what Grade?

OpenStudy (juan1857):

yeah, I was taught using arctan, but I think your way would also work

OpenStudy (kittiwitti1):

what class grade, or what grade I have in the course? o_o

OpenStudy (juan1857):

yeah

OpenStudy (juan1857):

class grade

OpenStudy (kittiwitti1):

College freshman, second semester.

OpenStudy (juan1857):

no wonder it seems like you know everything, lol

OpenStudy (kittiwitti1):

I do not know everything... xD but thank you for the compliment :P

OpenStudy (juan1857):

ok so back to the question, whats next lol

OpenStudy (kittiwitti1):

*brain died*

OpenStudy (kittiwitti1):

WAIT NO

OpenStudy (juan1857):

XD

OpenStudy (juan1857):

its ok, I'm assuming you got your own test to prepare for, Thanks

OpenStudy (kittiwitti1):

Okay so hahrom was right... \(3x-4y-10=0\rightarrow3r\cos{\theta}-4r\sin{\theta}=10\) Then you can factor r out of the left side: \(r(3\cos{\theta}-4\sin{\theta})=10\)

OpenStudy (juan1857):

ok, I'm having trouble getting \[\phi\]

OpenStudy (kittiwitti1):

then to get r you divide the other stuff from 10: \(r=\frac{10}{3\cos{\theta}-4\sin{\theta}}\)

OpenStudy (kittiwitti1):

Uh... I think to get \(\theta\) you would need to get the \(\arctan{\frac{y}{x}}\)? Arctangent is also \(\tan^{-1}{\theta}\)

OpenStudy (kittiwitti1):

HOLD ON I NEED TO ANSWER MY OWN QUESTION @ - @

OpenStudy (juan1857):

ok

OpenStudy (juan1857):

i used \[xcos \theta+ysin \theta-p=0\]

OpenStudy (kittiwitti1):

@Noobsteriscastic

OpenStudy (noobsteriscastic):

Is it solved?

OpenStudy (juan1857):

nope

OpenStudy (kittiwitti1):

He needs help finding the angle \(\theta\) but I have no idea how to do that, my brain is fried from my own polar conversions lol @Noobsteriscastic

OpenStudy (juan1857):

yeah what she said

OpenStudy (juan1857):

can you help @Noobsteriscastic?

OpenStudy (noobsteriscastic):

ok so you have 3x-4y=20 use x=rcos(theta), y=rsin(theta)

OpenStudy (juan1857):

ok

OpenStudy (juan1857):

I'm just gonna call it a day, Thanks @Noobsteriscastic ,@bahrom7893, and @kittiwitti1 once again

OpenStudy (noobsteriscastic):

You need to convert (x,y) to (r, theta) first of all find x, find y. To find x put y=0. To find y, put x=0. then use \[x^2 + y^2 =r^2 \] to find r and theta=arctan(y/x) to find find theta.

OpenStudy (juan1857):

sorry I just can't think at the moment Thanks anyways, I appreciate it

OpenStudy (noobsteriscastic):

ok well. how about you take a break and come back later?

OpenStudy (juan1857):

the problem is that once I take that break, I'll knock out 'till tomorrow, lol

OpenStudy (noobsteriscastic):

lol dude, you can't learn if you don't want to.

OpenStudy (juan1857):

ik, XD

OpenStudy (juan1857):

Thanks anyway @Noobsteriscastic